# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v4

@@ 4.3 The future of AI technologies
LLMs excel in applications involving complex tasks, advanced reasoning, data analysis, and understanding context. However, these models typically require high computational resources or cloud computing for development, operation and fine-tuning. A new trend emerging alongside LLMs is the development of Small language models (SLMs), such as Phi-3 by Microsoft [^61] https://azure.microsoft.com/en-us/blog/introducing-phi-3-redefining-whats-possible-with-slms/. SLMs offer promising solutions for regulated industries and sectors encountering scenarios where high-quality results are essential while keeping data ’on-site’. Their potential is particularly relevant when deploying more capable SLMs on smartphones and other mobile devices, allowing them to operate ‘at the edge’ without relying on cloud connectivity. Recent highly successful platforms, such as DeepSeek-V3 [420] and Qwen2.5-Max [421], are based on Mixture-of-Experts (MoE) models, which tackle complex problems by dividing them into simpler sub-tasks, each handled by a specialized “expert.”
@@
Finally, as stated in [2], the rapid advancement of AI technologies has revolutionized cultural experiences, often referred to as ‘CreaTech’—the convergence of the creative and digital sectors [422]. Such innovations not only reshape how people engage with art and creative work (e.g., through AR/VR/MR) but also drive the evolution of the technologies themselves.

### v5

@@ 4.3 The future of AI technologies
LLMs excel in applications involving complex tasks, advanced reasoning, data analysis, and understanding context. However, these models typically require high computational resources or cloud computing for development, operation and fine-tuning. A new trend emerging alongside LLMs is the development of Small language models (SLMs), such as Phi-3 by Microsoft [^61] https://azure.microsoft.com/en-us/blog/introducing-phi-3-redefining-whats-possible-with-slms/. SLMs offer promising solutions for regulated industries and sectors encountering scenarios where high-quality results are essential while keeping data ’on-site’. Their potential is particularly relevant when deploying more capable SLMs on smartphones and other mobile devices, allowing them to operate ‘at the edge’ without relying on cloud connectivity. Recent highly successful platforms, such as DeepSeek-V3 [deepseekv3] and Qwen2.5-Max [qwen25], are based on Mixture-of-Experts (MoE) models, which tackle complex problems by dividing them into simpler sub-tasks, each handled by a specialized “expert.”
@@
Finally, as stated in [Jeary2024], the rapid advancement of AI technologies has revolutionized cultural experiences, often referred to as ‘CreaTech’—the convergence of the creative and digital sectors [CreativeIndustriesCouncil2021]. Such innovations not only reshape how people engage with art and creative work (e.g., through AR/VR/MR) but also drive the evolution of the technologies themselves.
[…]
Research funding
[…]
Data Availability No datasets were generated or analysed during the current study.
[…]
Author Contributions Statement N.A. wrote the main manuscript text in section 1, 2, 3.1-3.5,3.8, 4, prepared all figures. F.Z. wrote the main manuscript text in section 3.6-3.7. D.B. wrote the main manuscript text in section 4. All authors reviewed the manuscript.